OpenStudy (anonymous):

Hey guy's :) I need some help again come take a look please? John began his job making $20 the first day. After that he was paid $6.00 per hour. Y = 6x + 20. Using the graph you drew for the equation, predict the pay for John if he works 7 hours after the first day.

OpenStudy (anonymous):

From my previous post of drawing linear line, Find x = 7 ( 7 hrs) => y = the pay! Double check y = f (7) = 6 * 7 + 20 = $62

OpenStudy (ash2326):

It's given John gets $20 + 6 per hour Let Y be the total earnings x= no. of hours worked \[Y=6x+20\] it's given he works 7 hours, so x=7 We get \[Y=6 \times 7+20=$62\] We can also find this from the graph of the line Let's plot it , we need two points put x=0, we get y=20 so one point is (0, 20) let's put y=0 so we get x=-20/6=-3.333 so other point is (-3.33, 0)  from the graph also we see y=$62